OpenStudy (dukelaxfan):

The point (−3, 1) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Make sure to show all work.

OpenStudy (dukelaxfan):

@mathstudent55 @agent0smith

OpenStudy (dukelaxfan):

@jhonyy9

OpenStudy (dukelaxfan):

@Mehek14 @Michele_Laino @Jaynator495

OpenStudy (dukelaxfan):

I know the first step it to find the radius

jhonyy9 (jhonyy9):

how you think it ? first you need drawing this point in the coordinates plane and after drawing this angle

OpenStudy (dukelaxfan):

Never mind, I figured it out

jhonyy9 (jhonyy9):

can you post it here ?

OpenStudy (dukelaxfan):

Find the radius \[r=\sqrt{-3^2+1^2}\]

OpenStudy (dukelaxfan):

which equals approx 3.2

OpenStudy (dukelaxfan):

Then just solve for each one

OpenStudy (dukelaxfan):

sin=y/r cos=x/r and tan=y/x

OpenStudy (dukelaxfan):

are we done here?

jhonyy9 (jhonyy9):

yes i think you are right

OpenStudy (dukelaxfan):

ok thx

jhonyy9 (jhonyy9):

np bye